a student missed 5 word problems on an test and received a grade of 62% If all the problems were of equal value, how many problems were on the test?
Answer by lynnlo(4176)

EXAMPLE:
ERIN MISSED 5 QUESTIONS ON A TEST,HER AVERAGE IS 83%
LET X=======NUMBER OF PROBLEMS ON TEST
(X-5)/X=.83=====================5 IS THE NUMBER SHE GOT WRONG,83 AVERAGE
MULTIPLY BOTH SIDES BY X
X-5=.83X
X-.83X=5
.17X=5
X=5/7
X=29.4~29 PROBLEMS WERE ON THE TEST